OpenStudy (anonymous):

If a triangle has a height of 12 in. and a base of 5 in., what is its area? A. 45 sq. in. B. 30 sq. in. C. 60 sq. in. D. 90 sq. in.

OpenStudy (calculusxy):

area of triangle=\[\frac{ 1 }{ 2 } (base \times height) \]

OpenStudy (calculusxy):

substitute. base is 5 and height is 12. so 12*5=60. half (0.50) of 60=30 (0.5*60)

OpenStudy (calculusxy):

answer. b

OpenStudy (anonymous):

thnk you